Find the equation of the tangent line at the given point on the following curve.

x^2+y^2=41, (-5,4)
1
Expert's answer
2020-04-14T16:14:32-0400

"2x+2yy'=0" hence "y'=-\frac{2x}{2y}=-\frac{x}{y}" then "y'=5\/4" if "(-5,4)" , hence equation of the tangent is "y-4=\frac{5}{4}(x+5)" or "5x-4y+41=0"